Samson inc will receive EUR 1 000 000 in one year. The interest rate on the euro is 3% p.aa and on the US dollar is 5%. The current spot rate for the euro is $1.00. If Samson uses a money market, how much will the hedge cost approximately?
Answer #1

The cost of the hedge has to be determined using Covered Interest Rate Parity.

The formula for Covered Interest Rate Parity is as follows:-

(1 + Interest Rate of Base Currency) = (Forward Foreign Exchange Rate / Current Spot Exchange Rate) * (1 + Interest Rate of Quoted Currency)

In the above problem, Euro is the base currency and Dollar is the Quoted currency

Putting the values:-

(1 + 0.03) = (Forward Foreign Exchange Rate / 1.00) * (1 + 0.05)

1.03 / 1.05 =(Forward Foreign Exchange Rate / 1.00)

Forward Foreign Exchange Rate = $ 0.9809 / Euro

Samson Inc will receive Euro 1,000,000 in one year

If it wants to hedge itself, it will have to sell Euro at the rate of $ 0.9809 / Euro, thereby getting = $ 980,900

At the prevailing Spot rate of $ 1 / Euro, if we sell Euro 1,000,000 we will get $ 1,000,000.

Hence, cost of hedging = (1,000,000 - 980,900) = $ 19,100